# Bidirectional Temporal Dynamics Modeling for EEG-based Driving Fatigue Recognition

### v1

@@ Abstract
I Introduction
@@
The main contributions of this work are summarized as follows
@@ II Related Works
II-A Implicit Temporal Aggregation
@@
II-B Multi-Scale Symmetric Temporal Modeling
@@
II-C Global Temporal Dependency Modeling
@@
II-D Lightweight and Physiologically Constrained Temporal Modeling
@@
III Methods
In this section, we propose an outstanding machine learning model that excels in driving fatigue detection on raw, unprocessed EEG signals. The overview of our model, DeltaGateNet, is shown in Fig. 1. In particular, the model consists of three simple but effective modules. Firstly, the Bidirectional Delta consists of no learnable parameters but is able to assist simple models to learn deeper representations of the EEG signal, which will be verified in Section V: Experimental Results. Secondly, the signals are passed to a Gated Temporal Convolution module, where EEG signals are further extracted and refined, turning into temporal encoded vectors of dimension . Finally, the EEG signals are passed into a multilayer perceptron for processing and classification, producing three fatigue prediction heads corresponding to the three classes.
III-A Bidirectional Delta Representations
@@
This approximates the time differential of EEG signals, which measures the discrete difference. This highlights rapid changes like a high-pass filter, removing slow drifts and offsets that are prevalent in unprocessed and unfiltered EEG signals. Then, the time derivatives are passed through the rectified linear unit that effectively separates positive differentials from negative differentials consisting of two tensors of shape where denotes the batch size, denotes the number of EEG input channels and denotes the number of timesteps. The rectified linear unit can be described by the equation:
@@
III-B Gated Temporal Convolution
@@
III-C Multilayer Perceptron
In this module, the pooled features of size are flattened to form a tensor of size . The two-dimensional tensor is fed into a multilayer perceptron that consists of three sequential blocks, each containing a linear, BatchNorm, LeakyReLU, and Dropout operation to produce three logits that correspond to the three classes. The detailed implementation can be visualized in Fig. 4. We then employ to convert logits to probabilities. The Softmax operation is denoted as:
@@ III-C Multilayer Perceptron
IV Experimental Design
IV-A Datasets
@@ IV-A1 SEED-VIG
IV-A2 Sustained Attention Driving Task (SADT)
@@
IV-B Baselines
For a comprehensive analysis of our model’s superior performance and reproducibility, we compared our model’s performance with five benchmark models, spanning from time-series classification models that aim to represent sequential data, deep learning CNN models that capture deep representations of EEG data, and EEG-based models that are tuned specifically for EEG data.
IV-B1 Time Series Classification Models
@@
IV-B2 Deep Learning CNN and Transformer-Based Models
@@
IV-B3 EEG-Based Models
EEGNet [44] employs depthwise and separable convolutions to construct a compact and versatile network architecture, which is widely recognized as the state-of-the-art model for diverse EEG classification tasks.
IV-C Implementation Details and Evaluation Metrics
@@ V Experimental Results
V-A Comparison Among Previous Models
V-A1 Intra-Subject Experiments
Table I presents the intra-subject classification performance, evaluated through 5-fold cross-validation, across the SEED-VIG and two SADT datasets for all compared models. The proposed DeltaGateNet consistently achieves the highest performance across all four evaluation metrics on every dataset.
On the SEED-VIG dataset, DeltaGateNet attains an Accuracy of , surpassing the next best method (Resnet_Wang, ) by nearly 4 percentage points. It also shows superior gains in Precision ( vs. ), Recall ( vs. ), and F1-score ( vs. ). Moreover, DeltaGateNet achieves dominating performance over models designed to capture EEG signals. DeltaGateNet achieves an eight percent gain over EEGNet on accuracy.
On the SADT 2022 dataset and the SADT 2952 dataset, DeltaGateNet once again surpasses all SOTA models. Surprisingly, time series classification models outperform both EEG-based models and transformer-based models on multiple datasets.
@@ V-A1 Intra-Subject Experiments
V-A2 Inter-Subject Experiments
Table II presents the more challenging inter-subject classification results, evaluated through 5-fold cross-subject validation, across the SEED-VIG and SADT datasets. In this rigorous cross-subject paradigm, our proposed DeltaGateNet demonstrates superior generalization by consistently achieving the highest accuracy on all three datasets.
@@ V-A2 Inter-Subject Experiments
V-B Ablation Studies
| Intrasubject Evaluation | ||||
| Model | Acc | Prec | Rec | F1 |
| MLP | ||||
| BidirectionalDelta + MLP | ||||
| GatedTemporalConv + MLP | ||||
| DeltaGateNet | ||||
| Intersubject Evaluation | ||||
| MLP | ||||
| Bidirectional Delta + MLP | ||||
| GatedTemporalConv + MLP | ||||
| DeltaGateNet |
| Intrasubject Evaluation | ||||
| Model | Acc | Prec | Rec | F1 |
| MLP | ||||
| BidirectionalDelta + MLP | ||||
| GatedTemporalConv + MLP | ||||
| DeltaGateNet | ||||
| Intersubject Evaluation | ||||
| MLP | ||||
| Bidirectional Delta + MLP | ||||
| GatedTemporalConv + MLP | ||||
| DeltaGateNet |
| Intrasubject Evaluation | ||||
| Model | Acc | Prec | Rec | F1 |
| MLP | ||||
| BidirectionalDelta + MLP | ||||
| GatedTemporalConv + MLP | ||||
| DeltaGateNet | ||||
| Intersubject Evaluation | ||||
| MLP | ||||
| Bidirectional Delta + MLP | ||||
| GatedTemporalConv + MLP | ||||
| DeltaGateNet |
V-B1 Influence of Each Individual Module
@@
In this section, we conduct a thorough investigation of the learning capabilities of our modules. We conducted rigorous experimentation and analysis of our model. Our results can be summarized in Tables III, IV, V. Furthermore, we evaluate our model on a more challenging inter-subject classification experiment to test our model’s generalization across different subjects. As shown in the tables, inter-subject performance is consistently lower than intra-subject performance across all models and datasets, reflecting the difficulty of cross-subject generalization. Despite this, our final model achieves the best performance in all inter-subject settings, demonstrating strong generalization capability.
@@
On Table III, in the intrasubject setting, using only an MLP yields poor performance . Adding the Bidirectional Delta module improves the accuracy to . The Gated Temporal Convolution + MLP achieves , while the full DeltaGateNet reaches , demonstrating the synergistic effect of both modules. This proves that our modules are capable of extracting features from different latent spaces. The Bidirectional Delta Module is able to improve the model’s accuracy by 11% demonstrating the robust capabilities of this module. Not only does it improve the performance of simple models such as MLP, but it also consists of no learnable parameters, leading to a more efficient design. On intersubject settings, our model improves Accuracy by 9.72% over the next best baseline (Bidirectional Delta + MLP), demonstrating similar learning capabilities on intrasubject settings.
@@
On SADT 2022 IV, in the intrasubject setting. the full model achieves Accuracy, compared to for Gated Temporal Conv + MLP and for Bidirectional Delta + MLP. This demonstrates its superior performance on bi-classification datasets, outperforming many SOTA models despite its simple design. In intersubject settings, the full model achieves Accuracy, compared to for Gated Temporal Conv + MLP and for Bidirectional Delta + MLP, showcasing its cross-domain capabilities on bi-classification datasets and its readiness for real-world driving scenarios.
@@
On SADT 2952 V, the full model again shows the best performance, confirming that both Bidirectional temporal modeling and channel-wise encoding are essential for capturing fatigue-related dynamics. DeltaGateNet attains Accuracy, outperforming Gated Temporal Conv + MLP and Bidirectional Delta + MLP on intersubject settings. Its performance is even more pronounced in the intrasubject setting, where its accuracy reaches an astonishing 96.84%.
@@
V-B2 Ablation Studies for Parameter Analysis
| Intrasubject Evaluation | ||||
| Kernel Size | Acc | Prec | Rec | F1 |
| K = 1 | ||||
| K = 3 | ||||
| K = 5 | ||||
| K = 7 | ||||
| Intersubject Evaluation | ||||
| Kernel Size | Acc | Prec | Rec | F1 |
| K = 1 | ||||
| K = 3 | ||||
| K = 5 | ||||
| K = 7 |
| Intrasubject Evaluation | ||||
| Step Size | Acc | Prec | Rec | F1 |
| S = 1 | ||||
| S = 2 | ||||
| S = 3 | ||||
| Intersubject Evaluation | ||||
| Step Size | Acc | Prec | Rec | F1 |
| S = 1 | ||||
| S = 2 | ||||
| S = 3 |
| Intrasubject Evaluation | ||||
| Kernel Size | Acc | Prec | Rec | F1 |
| K = 1 | ||||
| K = 3 | ||||
| K = 5 | ||||
| K = 7 | ||||
| Intersubject Evaluation | ||||
| Kernel Size | Acc | Prec | Rec | F1 |
| K = 1 | ||||
| K = 3 | ||||
| K = 5 | ||||
| K = 7 |
| Intrasubject Evaluation | ||||
| Step Size | Acc | Prec | Rec | F1 |
| S = 1 | ||||
| S = 2 | ||||
| S = 3 | ||||
| Intersubject Evaluation | ||||
| Step Size | Acc | Prec | Rec | F1 |
| S = 1 | ||||
| S = 2 | ||||
| S = 3 |
| Intrasubject Evaluation | ||||
| Kernel Size | Acc | Prec | Rec | F1 |
| K = 1 | ||||
| K = 3 | ||||
| K = 5 | ||||
| K = 7 | ||||
| Intersubject Evaluation | ||||
| Kernel Size | Acc | Prec | Rec | F1 |
| K = 1 | ||||
| K = 3 | ||||
| K = 5 | ||||
| K = 7 |
| Intrasubject Evaluation | ||||
| Step Size | Acc | Prec | Rec | F1 |
| S = 1 | ||||
| S = 2 | ||||
| S = 3 | ||||
| Intersubject Evaluation | ||||
| Step Size | Acc | Prec | Rec | F1 |
| S = 1 | ||||
| S = 2 | ||||
| S = 3 |
@@ V-B2 Ablation Studies for Parameter Analysis
Turning to tables VI and VII for the SEED-VIG dataset, we can see that parameters increased 10% on intrasubject evaluation when kernel size started from and ended at . Similar effects are shown on the intersubject evaluation paradigm as well. As for the step size, we are shown that the accuracy decreases with larger step sizes. We believe that this is because the approximation for the differential becomes less robust as the step size increases.
Tables VIII and IX show our parameter analysis for the SADT 2022 dataset. We can see that as the parameters approach our optimal parameters, not only does the accuracy increase gradually, but the variance also decreases concurrently, suggesting higher stability. The parameter analysis for the SADT 2952 dataset X XI concurs with the fact that our parameters are optimal.
The results confirm that the configuration and yield superior performance across both intrasubject and intersubject evaluation paradigms. The temporal convolution captures broad textual patterns while the fine-grained delta module extracts critical short-term Bidirectional changes. Their fused representation provides a robust and comprehensive feature space for vigilance estimation.
VI Conclusion
@@
Acknowledgments

### v2

@@ Abstract
Driving fatigue recognition, Electroencephalography (EEG), DeltaGateNet, Bidirectional Delta, Gated Temporal Convolution
1 Introduction
@@
The main contributions of this work are summarized as follows:
@@ 2 Related Works
2.1 Implicit Temporal Aggregation
@@
2.2 Multi-Scale Symmetric Temporal Modeling
@@
2.3 Global Temporal Dependency Modeling
@@
2.4 Lightweight and Physiologically Constrained Temporal Modeling
@@
3 Methods
In this section, we propose a novel machine learning model that excels in driving fatigue detection on raw, unprocessed EEG signals. The overview of our model, DeltaGateNet, is shown in Fig. 1. In particular, the model consists of three modules. Firstly, the Bidirectional Delta consists of no learnable parameters but is able to assist simple models to learn deeper representations of the EEG signal, which will be verified in Section V: Experimental Results. Secondly, the signals are passed to a Gated Temporal Convolution module, where EEG signals are further extracted and refined, turning into temporal encoded vectors of dimension . Finally, the EEG signals are passed into a multilayer perceptron for processing and classification, producing fatigue prediction heads corresponding to the number of classes.
3.1 Bidirectional Delta Representations
@@
where denotes the number of timesteps in the input. This approximates the time differential of EEG signals, which measures the discrete difference. This highlights rapid changes like a high-pass filter, removing slow drifts and offsets that are prevalent in unprocessed and unfiltered EEG signals. Then, the time derivatives are passed through the rectified linear unit that effectively separates positive differentials from negative differentials consisting of two tensors of shape where denotes the batch size, denotes the number of EEG input channels and denotes the number of timesteps. The rectified linear unit can be described by the equation:
@@
3.2 Gated Temporal Convolution
@@
3.3 Multilayer Perceptron
In this module, the pooled features of size are flattened to form a tensor of size . The two-dimensional tensor is fed into a multilayer perceptron that consists of three sequential blocks, each containing a linear, BatchNorm, LeakyReLU, and Dropout operation to produce logits that correspond to classes. We then employ to convert logits to probabilities. The Softmax operation is denoted as:
@@ 3.3 Multilayer Perceptron
4 Experimental Design
4.1 Datasets
@@ 4.1.1 SEED-VIG
4.1.2 Sustained Attention Driving Task (SADT)
@@
4.2 Baselines
For a comprehensive analysis of our model’s performance, we compared our model’s performance with eight benchmark models, spanning from time-series classification models that aim to represent sequential data, deep learning CNN models that capture deep representations of EEG data, and EEG-based models that are tuned specifically for EEG data.
4.2.1 Time Series Classification Models
@@
4.2.2 Deep Learning CNN and Transformer-Based Models
@@
4.2.3 EEG-Based Models
EEGNet [44] employs depthwise and separable convolutions to construct a compact and versatile network architecture, which is widely recognized as the state-of-the-art model for diverse EEG classification tasks. Modified TSception [45] enhances spatiotemporal feature extraction through a five-layer hierarchical temporal refinement strategy and Adaptive Average Pooling (ADP), offering improved stability for driver drowsiness and mental workload assessment. Additionally, GAT-CNN [46] integrates Graph Attention Networks with CNNs, utilizing a mutual information-driven adjacency matrix to model inter-channel dependencies for end-to-end driver fatigue detection.
4.3 Implementation Details and Evaluation Metrics
@@ 5 Experimental Results
5.1 Comparison Among Previous Models
5.1.1 Intra-Subject Experiments
Table 1 presents the intra-subject classification performance, evaluated through 5-fold cross-validation, across the SEED-VIG and two SADT datasets for all compared models. The proposed DeltaGateNet consistently achieves the highest performance across all four evaluation metrics on every dataset.
On the SEED-VIG dataset, DeltaGateNet attains an Accuracy of , surpassing the next best method (TSception, ) by nearly 3 percentage points. It also shows superior gains in Precision ( vs. ), Recall ( vs. ), and F1-score ( vs. ). Moreover, DeltaGateNet achieves an eight percent gain over EEGNet on accuracy.
On the SADT 2022 dataset and the SADT 2952 dataset, DeltaGateNet surpasses all SOTA models. Surprisingly, time series classification models outperform both EEG-based models and transformer-based models on multiple datasets.
@@ 5.1.1 Intra-Subject Experiments
5.1.2 Inter-Subject Experiments
Table 2 presents the more challenging inter-subject classification results, evaluated through 5-fold cross-subject validation, across the SEED-VIG and SADT datasets. In this rigorous cross-subject paradigm, our proposed DeltaGateNet demonstrates superior generalization by consistently achieving the highest accuracy on all three datasets.
@@ 5.1.2 Inter-Subject Experiments
5.2 Ablation Studies
| Model | SEED-VIG | SADT 2022 | SADT 2952 | |||||||||
| Acc | Prec | Rec | F1 | Acc | Prec | Rec | F1 | Acc | Prec | Rec | F1 | |
| Intrasubject Evaluation | ||||||||||||
| MLP | ||||||||||||
| Bidirectional Delta + MLP | ||||||||||||
| GatedTemporalConv + MLP | ||||||||||||
| DeltaGateNet | ||||||||||||
| Intersubject Evaluation | ||||||||||||
| MLP | ||||||||||||
| Bidirectional Delta + MLP | ||||||||||||
| GatedTemporalConv + MLP | ||||||||||||
| DeltaGateNet | ||||||||||||
5.2.1 Influence of Each Individual Module
@@
In this section, we conduct a thorough investigation of the learning capabilities of our modules. We conducted rigorous experimentation and analysis of our model. Our results can be summarized in Table 3. Furthermore, we evaluate our model on a more challenging inter-subject classification experiment to test our model’s generalization across different subjects. As shown in the tables, inter-subject performance is consistently lower than intra-subject performance across all models and datasets, reflecting the difficulty of cross-subject generalization. Despite this, our final model achieves the best performance in all inter-subject settings, demonstrating strong generalization capability.
@@
On SEED VIG 3, in the intrasubject setting, using only an MLP yields poor performance . Adding the Bidirectional Delta module improves the accuracy to . The Gated Temporal Convolution + MLP achieves , while the full DeltaGateNet reaches , demonstrating the synergistic effect of both modules. This proves that our modules are capable of extracting features from different latent spaces. The Bidirectional Delta Module is able to improve the model’s accuracy by 11% demonstrating the robust capabilities of this module. Not only does it improve the performance of simple models such as MLP, but it also consists of no learnable parameters, leading to a more efficient design. On intersubject settings, our model improves Accuracy by 9.72% over the next best baseline (Bidirectional Delta + MLP), demonstrating similar learning capabilities on intrasubject settings.
@@
On SADT 2022 3, in the intrasubject setting. the full model achieves Accuracy, compared to for Gated Temporal Conv + MLP and for Bidirectional Delta + MLP. In intersubject settings, the full model achieves Accuracy, compared to for Gated Temporal Conv + MLP and for Bidirectional Delta + MLP, showcasing its cross-domain capabilities on bi-classification datasets and its readiness for real-world driving scenarios.
@@
On SADT 2952 3, the full model again shows the best performance, confirming that both Bidirectional temporal modeling and channel-wise encoding are essential for capturing fatigue-related dynamics. DeltaGateNet attains Accuracy, outperforming Gated Temporal Conv + MLP and Bidirectional Delta + MLP on intersubject settings. Its performance is even more pronounced in the intrasubject setting, where its accuracy reaches an astonishing 96.84%.
@@
5.2.2 Ablation Studies for Parameter Analysis
| Kernel Size | SEED-VIG | SADT 2022 | SADT 2952 | |||||||||
| Acc | Prec | Rec | F1 | Acc | Prec | Rec | F1 | Acc | Prec | Rec | F1 | |
| Intrasubject Evaluation | ||||||||||||
| K = 1 | ||||||||||||
| K = 3 | ||||||||||||
| K = 5 | ||||||||||||
| K = 7 | ||||||||||||
| Intersubject Evaluation | ||||||||||||
| K = 1 | ||||||||||||
| K = 3 | ||||||||||||
| K = 5 | ||||||||||||
| K = 7 | ||||||||||||
| Step Size | SEED-VIG | SADT 2022 | SADT 2952 | |||||||||
| Acc | Prec | Rec | F1 | Acc | Prec | Rec | F1 | Acc | Prec | Rec | F1 | |
| Intrasubject Evaluation | ||||||||||||
| S = 1 | ||||||||||||
| S = 2 | ||||||||||||
| S = 3 | ||||||||||||
| Intersubject Evaluation | ||||||||||||
| S = 1 | ||||||||||||
| S = 2 | ||||||||||||
| S = 3 | ||||||||||||
@@ 5.2.2 Ablation Studies for Parameter Analysis
Turning to Table 4 for the SEED-VIG dataset, we can see that parameters increased 10% on intrasubject evaluation when kernel size started from and ended at . Similar effects are shown on the intersubject evaluation paradigm as well. As for the step size, we are shown that the accuracy decreases with larger step sizes. We believe that this is because the approximation for the differential becomes less robust as the step size increases.
Table 4 shows our parameter analysis for the SADT 2022 dataset. We can see that as the parameters approach our optimal parameters, not only does the accuracy increase gradually, but the variance also decreases concurrently, suggesting higher stability. The parameter analysis for the SADT 2952 dataset on Table 5 concurs with the fact that our parameters are optimal.
The results confirm that the configuration and yield better performance across both intrasubject and intersubject evaluation paradigms. The temporal convolution captures broad textual patterns while the fine-grained delta module extracts critical short-term Bidirectional changes. Their fused representation provides a robust and comprehensive feature space for vigilance estimation.
5.2.3 Frequency-Band Analysis of BidirectionalDelta Features
We evaluated the association between BidirectionalDelta energy and PERCLOS across canonical EEG frequency bands following rigorous preprocessing (bandpass – Hz, notch Hz, artifact rejection). Theta-band (– Hz) BidirectionalDelta energy exhibited the strongest positive correlation with PERCLOS (Pearson , , , samples), consistent with established neurophysiology wherein theta oscillations increase during drowsy states. Theta oscillations (4–8 Hz) are well-established electrophysiological markers of drowsiness and sleep onset [47, 48]. Alpha-band features showed comparable performance (), while beta-band features demonstrated the expected negative correlation (), confirming that the BidirectionalDelta module captures genuine arousal-related EEG dynamics rather than artifacts. The details can be visualized in Figure 5. This theta-specific correlation pattern aligns precisely with known neurophysiology of vigilance decrements and provides convergent validity that BidirectionalDelta captures genuine drowsiness-related EEG dynamics rather than artifact. The module therefore offers a computationally efficient approach to extracting theta-band temporal features for real-time fatigue monitoring.
6 Conclusion
@@
Acknowledgment
We would like to thank Professor Xiaomin Ouyang, Professor Zhihong Li, and Professor Nevin L. Zhang for their invaluable guidance on our research paper.